OpenStudy (anonymous):

What is the slope of a line that passes through the point (−1, 1) and is parallel to a line that passes through (4, 6) and (−1, −4)?

OpenStudy (anonymous):

\[m = \frac{-4-6}{-1-4}=\frac{-10}{-5}=2 \]

OpenStudy (anonymous):

the point (-1,1) is needed to find the equation of the line. parallel lines have identical slopes. for perpendicular or normal lines, the product of their slopes is "-1"